# Efficient Uncertainty Propagation in Bayesian Two-Step Procedures

### v1

@@ Abstract
Keywords Bayesian Inference  $\cdot$ Importance Sampling  $\cdot$ Multiple Imputation  $\cdot$ Surrogate Modeling  $\cdot$ Uncertainty Propagation
@@ 1 Introduction
Many modeling approaches in science and engineering adopt a two-step framework, wherein an initial modeling or preprocessing phase is followed by a subsequent stage focused on inference or prediction [1, 2, 3, 4, 5]. While effective, such two-step procedures introduce additional complexity – particularly in the treatment of uncertainty. A key challenge in two-stage inference lies in accurately propagating uncertainty across steps to ensure the reliability and trustworthiness of the final outcomes [6, 7]. Addressing this challenge often entails computationally intensive and time-consuming procedures. In this paper, we propose a general methodology for efficient and robust uncertainty propagation within Bayesian two-step modeling frameworks. We demonstrate the effectiveness of our approach in two widely encountered statistical scenarios that demand careful uncertainty management, as described in the following sections.
A prominent example of a two-step modeling approach is the use of surrogate models in scenarios where simulations of complex phenomena are required – such as in the natural sciences [8, 9]. These simulations often involve solving differential equations that lack closed-form solutions and must be addressed through computationally expensive numerical methods [10]. To alleviate these computational challenges, surrogate models [11, 12] are employed as simplified approximations of the true system, enabling significantly faster evaluations while preserving essential behaviors of the original model. Surrogate modeling has found broad application across domains such as systems biology [13, 14], water resource management [15, 16], and climate modeling [17]. In a typical two-step surrogate modeling approach [18], the complex simulator is first approximated using a simpler surrogate trained on a limited set of outputs from the true model. This introduces uncertainty stemming from both limited training data and potential model mismatch. In the second step, inference is performed using the trained surrogate to estimate parameters based on observed data. To ensure reliable parameter estimates, it is critical to account for and appropriately propagate the uncertainties introduced during surrogate construction into the inference stage [19].
Another context in which a two-step procedure is essential arises when working with datasets containing missing values. Rather than discarding incomplete records – which may lead to biased results or information loss – imputation techniques are employed to estimate the missing entries [20]. Among these, Multivariate Imputation by Chained Equations (MICE) [21] stands out as a widely adopted method [22, 23, 24]. MICE addresses uncertainty from missingness by generating multiple complete datasets, each containing plausible imputations for the missing values [21]. In the second step, a statistical model is independently fitted to each imputed dataset, and the resulting posterior distributions are combined. While this approach effectively incorporates imputation uncertainty, it can become computationally demanding, especially when complex models are used. Nevertheless, it yields more robust and comprehensive posterior estimates by properly accounting for the uncertainty introduced during imputation.
In both examples – surrogate modeling and missing data handling – the key challenge is propagating the uncertainty from the first to the second step. A major drawback of existing methods is their high computational cost, as they require separate model fits for each realization of the uncertainty quantity (e.g. surrogate parameters or imputed datasets). This issue is particularly acute in the Bayesian setting, where posterior distributions are often approximated using costly sampling algorithms [25, 26], most notably Markov Chain Monte Carlo (MCMC) [27].
For example, if 100 realizations $\tau^{(1)},\dots,\tau^{(100)}$ are generated in the first step, this would typically necessitate estimating 100 separate models – each costly in terms of time and computational resources (see Figure 1(a)). To overcome this limitation, we propose an efficient approach that significantly reduces the computational overhead while preserving the trustworthiness of uncertainty propagation. Our approach leverages importance sampling techniques, allowing us to approximate the desired posterior distributions without running MCMC or comparably costly algorithms for each realization $\tau^{(i)}$. Specifically, we employ Pareto Smoothed Importance Sampling (PSIS) [28] and Importance Weighted Moment Matching (IWMM) [29]. With this strategy, MCMC is required only for a small subset of realizations of $\tau$, reducing computational demands. For the remaining realizations, we approximate the corresponding posterior distributions directly within the posterior space using advanced importance sampling (see Figure 1(b)). This approach offers a significant gain in efficiency while preserving the accuracy of the propagated uncertainty, making it a practical scalable solution for complex two-step modeling workflows.
@@ 2 Method
In this section, we present our proposed methodology for addressing uncertainty propagation within Bayesian two-step modeling frameworks. We begin by formally defining the problem setting and introducing the key components required for our approach. Section 2.3 then describes the core of our method – an iterative procedure designed for efficient and reliable uncertainty propagation across the two stages. Finally, we discuss two specific application scenarios that require additional notation and domain-specific considerations.
@@ 2.1 Problem Definition
[…]
@@
In many practical scenarios, the number of samples $m$ available is limited. Additionally, evaluating $p(\theta\mid\tau^{(i)})$ for each $i=1,\dots,m$ can be computationally expensive [32]. In the setting of Equation (2), the posterior distribution of $\theta$ given a $\tau^{(i)}$, $p(\theta\mid\tau^{(i)})$, cannot be accessed directly and must instead be approximated through expensive sampling algorithms [25], for example, via MCMC [26]. For each $\tau^{(i)}$, we then obtain $S$ samples $\{\theta^{(s)}_{i}\}_{s=1}^{S}\sim p(\theta\mid\tau^{(i)})$ from the corresponding posterior distribution. By treating all $m\cdot S$ draws as one large sample, we can use them to approximate the marginalized posterior distribution $p(\theta\mid C)$ [33].
@@
[…]
To ensure $\tau^{*}$ is representative of the entire sample, it can be chosen as a summary statistic, such as the mean or median of the $m$ samples, depending on the characteristics of $\tau$. While this approach is computationally efficient – requiring only a single MCMC run – it fails to propagate the uncertainty in $p(\tau\mid C)$, almost surely leading to overconfidence in the resulting marginalized posterior distribution $p(\theta\mid C)$.
A more accurate approximation of $p(\theta\mid C)$ requires using all available samples via the Monte Carlo estimate in Equation (2). To make this approach computationally feasible, the number of required MCMC runs needs to be reduced. To this end, we propose a combination of importance sampling techniques – specifically, Pareto smoothed importance sampling (PSIS) [28] and importance weighted moment matching (IWMM) [29], which are described in detail in Section 2.2. These methods allow us to approximate the distributions $p(\theta\mid\tau^{(i)})$ for $\{\tau^{(i)}\}_{i=1}^{m}$ within the posterior space as shown in Figure 1(b). This facilitates a more efficient calculation of the Monte Carlo estimate in Equation (2) than running costly sampling algorithms for each value of $\{\tau^{(i)}\}_{i=1}^{m}$ separately (see Figure 1(a) for comparison).
@@ 2.2 Method Components
Our objective is to approximate the distribution $p(\theta\mid\tau^{(i)})$ without relying on computationally expensive sampling algorithms. In the following, for simplicity, we refer only to MCMC to describe this class of expensive but highly accurate algorithms. Instead of running MCMC to access the distribution, we utilize a proposal distribution $p(\theta\mid\tau^{*})$ and apply Pareto smoothed importance sampling (PSIS). When PSIS alone is insufficient to accurately approximate $p(\theta\mid\tau^{(i)})$, we enhance the approximation using importance weighted moment matching (IWMM).
Importance Sampling The key idea behind importance sampling is to approximate properties of the target distribution $p(\theta\mid\tau^{(i)})$ without drawing samples directly from it. This allows, among others, to approximate expectations $\mathbb{E}_{p(\theta|\tau^{(i)})}(h(\theta)\mid\tau^{(i)})$ over the target for any function $h$ [30].
@@
[…]
@@
$\mathbb{E}_{p(\theta|\tau^{(i)})}(h(\theta)\mid\tau^{(i)})\approx\frac{\sum_{s=1}^{S}h(\theta^{(s)}_{*})r_{i}(\theta^{(s)}_{*})}{\sum_{s=1}^{S}r_{i}(\theta^{(s)}_{*})}=\sum_{s=1}^{S}h(\theta^{(s)}_{*})w_{i}(\theta^{(s)}_{*}),$ (5)
[…]
Here, $r_{i}(\theta^{(s)}_{*})$, for $s=1,\dots,S$, are referred to as importance ratios and $w_{i}(\theta^{(s)}_{*})$ are the corresponding importance weights. Using the same proposal distribution $p(\theta\mid\tau^{*})$, we can compute importance weights for each of the posterior distributions $p(\theta\mid\tau^{(i)})$ corresponding to the values $\tau^{(i)}\in\{\tau^{(i)}\}_{i=1}^{m}$. For importance sampling to be effective and yield reliable approximations, the target and proposal distributions must be sufficiently similar and share the same support. Since both $\tau^{(i)}$ and $\tau^{*}$ are drawn from the same distribution $p(\tau\mid C)$, the corresponding posteriors $p(\theta\mid\tau^{(i)})$ and $p(\theta\mid\tau^{*})$ typically share similar structure and support, which enables us to apply importance sampling effectively.
@@
Pareto Smoothed Importance Sampling Importance sampling, and consequently importance resampling, can be unreliable if the proposal distribution is not well-suited for the target distribution [34]. This issue can be diagnosed by examining the distribution of the importance weights $\{w_{i}(\theta^{(s)}_{*})\}_{s=1}^{S}$. If the distribution shows to have excessively heavy tails, importance sampling tends to fail, leading to unreliable results. To address this problem, various adaptions of importance sampling have been developed [35, 36]. In the following, we apply Pareto smoothed importance sampling (PSIS) [28], since it is a reliable and trustworthy variant of importance sampling that outperforms other methods in regard of stabilizing importance sampling estimates.
@@
If $\hat{k}_{i}<0.7$, the Pareto-smoothed importance weights $w_{i}$ are considered reliable and can be used to resample from $\{\theta^{(s)}_{*}\}_{s=1}^{S}$ to obtain $\{\theta^{(s)}_{i}\}_{s=1}^{S}\sim p(\theta\mid\tau^{(i)})$. However, when $\hat{k}_{i}\geq 0.7$, the proposal distribution is deemed insufficently close to the target distribution for reliable importance sampling. In this case, the importance weights exhibit heavy-tailed behavior, and the Pareto smoothing can no longer adequately control the estimator’s error characteristics. Specifically, the bias introduced by smoothing the extreme weights begins to dominate over the estimator’s variance, undermining both the accuracy and interpretability of the resulting estimates [28]. As a result, the PSIS-based approximation may be unstable or biased, and a more complex and advanced method might be necessary to successfully approximate the target distribution. To address this, we employ importance-weighted moment matching (IWMM), an adaptive importance sampling method described in detail below.
Importance Weighted Moment Matching The method introduced by Paananen et al. (2021) can be employed when PSIS fails, as indicated by a $\hat{k}>k_{\text{threshold}}$. Importance weighted moment matching (IWMM) is an adaptive importance sampling method, where the proposal distribution is iteratively updated to better approximate the target distribution. In general, these methods consist of three steps: (i) generating draws from the proposal distribution, (ii) computing the importance weights for these draws, and (iii) adapting the proposal distribution based on the computed weights [34].
In the specific case of IWMM, step (i) is omitted, meaning that no resampling occurs. Instead, the existing draws are transformed. In step (iii), the samples are transformed using an affine transformation. While these transformations can vary in complexity, for simplicity, we focus on basic affine transformations, since both the transformation and its Jacobian are computationally cheap [29]. To approximate the expectation $\mathbb{E}_{p(\theta\mid\tau^{(i)})}(h(\theta)\mid\tau^{(i)})$ with a set of draws $\{\theta^{(s)}_{*}\}_{s=1}^{S}$ from the proposal distribution $p(\theta\mid\tau^{*})$, we use a matrix $\mathbf{A}$ to represent a linear map and a vector $\mathbf{b}$ to represent a translation. A draw $\theta^{(s)}_{*}$ is then transformed as follows [29]:
@@
$\breve{\theta}^{(s)}_{*}=T_{1}(\theta^{(s)}_{*})=\theta^{(s)}_{*}-\bar{\theta}_{*}+\tilde{\theta}_{*}\text{ with }\bar{\theta}_{*}=\frac{1}{S}\sum_{s=1}^{S}\theta^{(s)}_{*}\text{ and }\tilde{\theta}_{*}=\frac{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})\theta^{(s)}_{*}}{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})}\,.$ (9)
@@
[…]
@@
[…]
@@
Although IWMM is computationally more expensive than PSIS, we employ it in cases where PSIS fails to provide reliable approximations of the target distribution. Figure 2 illustrates representative posterior distributions from two imputed datasets in the simulation study described in Section 3.1, within the special case of the missing data problem (see Section 2.4). The proposal distribution (shown in red) was obtained via MCMC, and the true target distribution (yellow) – which is typically unknown without running MCMC – serves as the reference. The PSIS-based approximation (light blue), constructed via importance resampling from the proposal, fails to accurately capture the shape and location of the target distribution. This poor approximation is flagged by a high Pareto shape parameter, $\hat{k}=0.93>k_{\mathrm{threshold}}$, indicating that the results are unreliable. In contrast, using additionally IWMM results in a substantially improved approximation (dark blue), with resampled draws that closely match the target distribution. This is confirmed by a much lower shape parameter, $\hat{k}_{\mathrm{MM}}=0.27<k_{\mathrm{threshold}}$, indicating a successful and trustworthy approximation.
@@
We then define the mixture proposal $p_{\rm mix}(\theta\mid\{\tau^{(j)}\})$ as
@@
$r_{i}(\theta^{(s)}_{*})$ $=\frac{p(\theta^{(s)}_{*}\mid\tau^{(i)})}{p_{\rm mix}(\theta_{*}^{(s)}\mid\{\tau^{(j)}\})}=\frac{p(\theta^{(s)}_{*}\mid\tau^{(i)})}{\frac{1}{\mid I_{\text{mix}}\mid}\sum_{j\in I_{\text{mix}}}p(\theta_{*}^{(s)}\mid\tau^{(j)})}$ (15)
@@
In the standard setting of self-normalized importance sampling, the normalizing constants – here the marginal likelihoods $p(\tau^{(i)})$ – can be ignored since the self-normalization automatically corrects for them. However, in the case of a mixture proposal, the constants of the mixture components cannot be factored out of the summation in the denominator, and thus self-normalization is no longer directly applicable. Therefore, to evaluate Equation (18), we require either the exact values or relative magnitudes of the marginal likelihoods $p(\tau^{(j)})$ for all $j\in I_{\text{mix}}$. For the purpose of estimating $p(\tau^{(j)})$, we use bridge sampling [39], a method for estimating normalizing constants in complex posterior distributions [40]. With these estimates, the mixture-based importance weights can be accurately computed, even though it adds computational complexity.
In our approach, we employ both single and mixture proposal distributions for PSIS, whereas IWMM is applied only with single proposals. Preliminary investigations (not shown here) indicated that using these mixtures within IWMM performs worse than using just a single posterior $p(\theta\mid\tau^{*})$ as proposal. We attribute this to the nature of the IWMM method, which relies on affine transformations to shift and scale the proposal distribution to match the moments of the weight distribution [29]. A single posterior distribution is more likely to share structural similarities – differing primarily in location and scale – with another posterior (the target) than a multi-modal mixture distribution. Based on these findings, we opted not to further consider mixture proposals when applying IWMM.

### v2

@@ Abstract
keywords:
Bayesian Inference, Importance Sampling, Multiple Imputation, Surrogate Modeling, Uncertainty Propagation
@@ 1 Introduction
Many modeling approaches in science and engineering adopt a two-step framework, wherein an initial modeling or preprocessing phase is followed by a subsequent stage focused on inference or prediction [46, 43, 39, 72, 17]. While effective, such two-step procedures introduce additional complexity — particularly in the treatment of uncertainty. A key challenge in two-stage inference lies in accurately propagating uncertainty across steps to ensure the reliability and trustworthiness of the final outcomes [73, 59]. Addressing this challenge often entails computationally intensive and time-consuming procedures. In this paper, we propose a general methodology for efficient and robust uncertainty propagation within Bayesian two-step modeling frameworks. We demonstrate the effectiveness of our approach in two widely encountered statistical scenarios that demand careful uncertainty management, as described in the following sections.
@@
A prominent example of a two-step modeling approach is the use of surrogate models in scenarios where simulations of complex phenomena are required — such as in the natural sciences [40, 21]. These simulations often involve solving differential equations that lack closed-form solutions and must be addressed through computationally expensive numerical methods [64]. To alleviate these computational challenges, surrogate models [83, 42] are employed as simplified approximations of the true system, enabling significantly faster evaluations while preserving essential behaviors of the original model. Surrogate modeling has found broad application across domains such as systems biology [67, 1], water resource management [65, 25], and climate modeling [41]. In a typical two-step surrogate modeling approach [66], the complex simulator is first approximated using a simpler surrogate trained on a limited set of outputs from the true model. This introduces uncertainty stemming from both limited training data and potential model mismatch. In the second step, inference is performed using the trained surrogate to estimate parameters based on observed data. To ensure reliable parameter estimates, it is critical to account for and appropriately propagate the uncertainties introduced during surrogate construction into the inference stage [63].
Another context in which a two-step procedure is essential arises when working with datasets containing missing values. Rather than discarding incomplete records — which may lead to biased results or information loss — imputation techniques are employed to estimate the missing entries [7]. Among these, Multivariate Imputation by Chained Equations (MICE) [6] stands out as a widely adopted method [70, 78, 71]. MICE addresses uncertainty from missingness by generating multiple complete datasets, each containing plausible imputations for the missing values [6]. In the second step, a statistical model is independently fitted to each imputed dataset, and the resulting posterior distributions are combined. While this approach effectively incorporates imputation uncertainty, it can become computationally demanding, especially when complex models are used. Nevertheless, it yields more robust and comprehensive posterior estimates by properly accounting for the uncertainty introduced during imputation.
In both examples — surrogate modeling and missing data handling — the key challenge is propagating the uncertainty from the first to the second step. A major drawback of existing methods is their high computational cost, as they require separate model fits for each realization of the uncertainty quantity (e.g. surrogate parameters or imputed datasets). This issue is particularly acute in the Bayesian setting, where posterior distributions are often approximated using costly sampling algorithms [26, 68], most notably Markov Chain Monte Carlo (MCMC) [4].
A closely related line of work arises in cut models (or cutting feedback), where inference is modularized to prevent information from a downstream (potentially misspecified) module from feeding back into an upstream module [2, 47, 61]. Therefore, cut models can be viewed as Bayesian two-step procedures and, in this sense, constitute a special case of our framework. In practice, most cut methods fall into two broad classes. First, two-step Monte Carlo (multiple-imputation style) approaches draw upstream quantities from $p(\phi\mid Z)$ and refit the downstream model for each draw [61]. Second, tailored algorithms approximate the cut distribution more directly, for example via tempered or modified transition kernels (motivated in part by the failure of naive cut samplers to target a well-defined stationary distribution) [61], stochastic-approximation methods such as SACut [45], or semi-modular approaches that downweight feedback rather than removing it entirely [13]. Beyond MCMC-based implementations, alternative strategies include posterior-bootstrap procedures for robust modular inference [62], generalized-Bayes formulations of cutting feedback [22], Gaussian-process emulation of conditional posteriors to reduce the number of costly refits [35], sequential Monte Carlo for cut posterior computation [49], and variational approaches that cast cutting feedback as constrained optimization [81]. Unbiased MCMC estimators based on couplings provide another general tool for estimating expectations under such targets [37]. While these approaches address important conceptual and computational issues, many still require repeated second-step MCMC runs or other substantial additional computation, reinforcing the need for efficient uncertainty propagation strategies in Bayesian two-step workflows.
For example, if 100 realizations $\tau^{(1)},\dots,\tau^{(100)}$ are generated in the first step, this would typically necessitate estimating 100 separate models — each costly in terms of time and computational resources (see Figure 1(a)). To overcome this limitation, we propose an efficient approach that significantly reduces the computational overhead while preserving the trustworthiness of uncertainty propagation. Our approach leverages importance sampling techniques, allowing us to approximate the desired posterior distributions without running MCMC or comparably costly algorithms for each realization $\tau^{(i)}$. Specifically, we employ Pareto Smoothed Importance Sampling (PSIS) [77] and Importance Weighted Moment Matching (IWMM) [56]. With this strategy, MCMC is required only for a small subset of realizations of $\tau$, reducing computational demands. For the remaining realizations, we approximate the corresponding posterior distributions directly within the posterior space using advanced importance sampling (see Figure 1(b)). This approach offers a significant gain in efficiency while preserving the accuracy of the propagated uncertainty, making it a practical scalable solution for complex two-step modeling workflows.
@@ 2 Method
In this section, we present our proposed methodology for addressing uncertainty propagation within Bayesian two-step modeling frameworks. We begin by formally defining the problem setting and introducing the key components required for our approach. Section 2.3 then describes the core of our method — an iterative procedure designed for efficient and reliable uncertainty propagation across the two stages. Finally, we discuss two specific application scenarios that require additional notation and domain-specific considerations.
@@ 2.1 Problem Definition
$\int p(\theta\mid\tau,C)p(\tau\mid C)\,d\tau\approx\frac{1}{m}\sum_{i=1}^{m}p\left(\theta\mid\tau^{(i)},C\right).$ (2)
@@
In many practical scenarios, the number of samples $m$ available is limited. For example in surrogate model applications, running the surrogate for many values of $\tau$ can be computationally demanding and therefore time- and cost-intensive. Similarly, in the multiple-imputation setting, increasing $m$ requires running the MICE algorithm many times, which can be expensive depending on the dataset size and missingness structure. Additionally, evaluating $p(\theta\mid\tau^{(i)})$ for each $i=1,\dots,m$ can be computationally expensive [51]. In the setting of Equation (2), the posterior distribution of $\theta$ given a $\tau^{(i)}$, $p(\theta\mid\tau^{(i)})$, cannot be accessed directly and must instead be approximated through expensive sampling algorithms [26], for example, via MCMC [68]. For each $\tau^{(i)}$, we then obtain $S$ samples $\{\theta^{(s)}_{i}\}_{s=1}^{S}\sim p(\theta\mid\tau^{(i)})$ from the corresponding posterior distribution. By treating all $m\cdot S$ draws as one large sample, we can use them to approximate the marginalized posterior distribution $p(\theta\mid C)$} [44].
@@
$p(\theta\mid C)=\int p(\theta\mid\tau)p(\tau\mid C)\,d\tau\approx p(\theta\mid\tau^{*})\,.$ (3)
To ensure $\tau^{*}$ is representative of the entire sample, it can be chosen as a summary statistic, such as the mean or median of the $m$ samples, depending on the characteristics of $\tau$. While this approach is computationally efficient — requiring only a single MCMC run — it fails to propagate the uncertainty in $p(\tau\mid C)$, almost surely leading to overconfidence in the resulting marginalized posterior distribution $p(\theta\mid C)$.
A more accurate approximation of $p(\theta\mid C)$ requires using all available samples via the Monte Carlo estimate in Equation (2). To make this approach computationally feasible, the number of required MCMC runs needs to be reduced. To this end, we propose a combination of importance sampling techniques — specifically, Pareto Smoothed Importance Sampling (PSIS) [77] and importance weighted moment matching (IWMM) [56], which are described in detail in Section 2.2. These methods allow us to approximate the distributions $p(\theta\mid\tau^{(i)})$} for $\{\tau^{(i)}\}_{i=1}^{m}$ within the posterior space as shown in Figure 1(b). This facilitates a more efficient calculation of the Monte Carlo estimate in Equation (2) than running costly sampling algorithms for each value of $\{\tau^{(i)}\}_{i=1}^{m}$ separately (see Figure 1(a) for comparison).
@@ 2.2 Method Components
Our objective is to approximate the distribution $p(\theta\mid\tau^{(i)})$ without relying on computationally expensive sampling algorithms. In the following, for simplicity, we refer only to MCMC to describe this class of expensive but highly accurate algorithms. Instead of running MCMC to access the distribution, we utilize a proposal distribution $p(\theta\mid\tau^{*})$ and apply PSIS. When PSIS alone is insufficient to accurately approximate $p(\theta\mid\tau^{(i)})$, we enhance the approximation using IWMM.
Importance Sampling The key idea behind importance sampling is to approximate properties of the target distribution $p(\theta\mid\tau^{(i)})$ without drawing samples directly from it. This allows, among other things, to approximate expectations $\mathbb{E}_{p(\theta|\tau^{(i)})}(h(\theta)\mid\tau^{(i)})$ over the target for any function $h$ [31].
@@
$\mathbb{E}_{p(\theta\mid\tau^{(i)})}(h(\theta)\mid\tau^{(i)})=\frac{\int h(\theta)p(\theta\mid\tau^{(i)})d\theta}{\int p(\theta\mid\tau^{(i)})d\theta}$ (4)
$=\frac{\int h(\theta)\frac{p(\theta\mid\tau^{(i)})}{p(\theta\mid\tau^{*})}p(\theta\mid\tau^{*})d\theta}{\int\frac{p(\theta\mid\tau^{(i)})}{p(\theta\mid\tau^{*})}p(\theta\mid\tau^{*})d\theta}\,.$ (5)
@@
$\mathbb{E}_{p(\theta|\tau^{(i)})}(h(\theta)\mid\tau^{(i)})$ $\approx\frac{\sum_{s=1}^{S}h(\theta^{(s)}_{*})r_{i}(\theta^{(s)}_{*})}{\sum_{s=1}^{S}r_{i}(\theta^{(s)}_{*})}$ (6)
$=\sum_{s=1}^{S}h(\theta^{(s)}_{*})w_{i}(\theta^{(s)}_{*}),$ (7)
$\text{ with }\;r_{i}(\theta^{(s)}_{*})$ $=\frac{p(\theta^{(s)}_{*}\mid\tau^{(i)})}{p(\theta^{(s)}_{*}\mid\tau^{*})}\;$ (8)
$\text{ and }\;w_{i}(\theta^{(s)}_{*})$ $=\frac{r_{i}(\theta^{(s)}_{*})}{\sum_{s=1}^{S}r_{i}(\theta^{(s)}_{*})}\,$, (9)
$s=1,\dots,S$. Here, $r_{i}(\theta^{(s)}_{*})$, for $s=1,\dots,S$, are referred to as importance ratios and $w_{i}(\theta^{(s)}_{*})$ are the corresponding importance weights. Using the same proposal distribution $p(\theta\mid\tau^{*})$, we can compute importance weights for each of the posterior distributions $p(\theta\mid\tau^{(i)})$ corresponding to the values $\tau^{(i)}\in\{\tau^{(i)}\}_{i=1}^{m}$. For importance sampling to be effective and yield reliable approximations, the target and proposal distributions must be sufficiently similar and share the same support. Since both $\tau^{(i)}$ and $\tau^{*}$ are drawn from the same distribution $p(\tau\mid C)$, the corresponding posteriors $p(\theta\mid\tau^{(i)})$ and $p(\theta\mid\tau^{*})$ typically share similar structure and support, which enables us to apply importance sampling effectively.
@@
Pareto Smoothed Importance Sampling Importance sampling, and consequently importance resampling, can be unreliable if the proposal distribution is not well-suited for the target distribution [5]. This issue can be diagnosed by examining the distribution of the importance weights $\{w_{i}(\theta^{(s)}_{*})\}_{s=1}^{S}$. If the distribution shows to have excessively heavy tails, importance sampling becomes unstable, often with very large or even infinite variance, leading to unreliable estimates [36, 77]. To address this problem, various adaptations of importance sampling have been developed [36, 52]. In the following, we apply PSIS [77], since it is a reliable and trustworthy variant of importance sampling that outperforms other methods in regard of stabilizing importance sampling estimates.
@@
If $\hat{k}_{i}<0.7$, the Pareto-smoothed importance weights $w_{i}$ are considered reliable and can be used to resample from $\{\theta^{(s)}_{*}\}_{s=1}^{S}$ to obtain $\{\theta^{(s)}_{i}\}_{s=1}^{S}\sim p(\theta\mid\tau^{(i)})$}. However, when $\hat{k}_{i}\geq 0.7$, the proposal distribution is deemed insufficiently close to the target distribution for reliable importance sampling. In this case, the importance weights exhibit heavy-tailed behavior, and the Pareto smoothing can no longer adequately control the estimator’s error characteristics. Specifically, the bias introduced by smoothing the extreme weights begins to dominate over the estimator’s variance, undermining both the accuracy and interpretability of the resulting estimates [77]. As a result, the PSIS-based approximation may be unstable or biased, and a more complex and advanced method might be necessary to successfully approximate the target distribution. To address this, we employ importance-weighted moment matching (IWMM), an adaptive importance sampling method described in detail below.
Importance Weighted Moment Matching The method introduced by [56] can be employed when PSIS fails, as indicated by a $\hat{k}>k_{\text{threshold}}$. Importance weighted moment matching (IWMM) is an adaptive importance sampling method, where the proposal distribution is iteratively updated to better approximate the target distribution. In general, these methods consist of three steps: (i) generating draws from the proposal distribution, (ii) computing the importance weights for these draws, and (iii) adapting the proposal distribution based on the computed weights [5].
In the specific case of IWMM, step (i) is omitted, so that the same samples are transformed directly and no resampling is necessary. With these samples, the importance weights are calculated in step (ii). In step (iii), the samples are transformed using an affine transformation. While these transformations can vary in complexity, for simplicity, we focus on basic affine transformations, since both the transformation and its Jacobian are computationally cheap [56]. To approximate the expectation $\mathbb{E}_{p(\theta\mid\tau^{(i)})}(h(\theta)\mid\tau^{(i)})$ with a set of draws $\{\theta^{(s)}_{*}\}_{s=1}^{S}$ from the proposal distribution $p(\theta\mid\tau^{*})$, we use a matrix $\mathbf{A}$ to represent a linear map and a vector $\mathbf{b}$ to represent a translation. A draw $\theta^{(s)}_{*}$ is then transformed as follows [56]:
@@
$\breve{\theta}^{(s)}_{*}$ $=T_{1}(\theta^{(s)}_{*})=\theta^{(s)}_{*}-\bar{\theta}_{*}+\tilde{\theta}_{*}$ (12)
$\text{ with }\bar{\theta}_{*}=\frac{1}{S}\sum_{s=1}^{S}\theta^{(s)}_{*}$ (13)
$\text{ and }\tilde{\theta}_{*}=\frac{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})\theta^{(s)}_{*}}{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})}\,.$ (14)
@@
 $\text{ with }\mathbf{v}=\frac{1}{S}\sum_{s=1}^{S}(\theta^{(s)}_{*}-\bar{\theta}_{*})\circ(\theta^{(s)}_{*}-\bar{\theta}_{*})$ (16)
 $\text{ and }\tilde{\mathbf{v}}=\frac{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})(\theta^{(s)}_{*}-\bar{\theta}_{*})\circ(\theta^{(s)}_{*}-\bar{\theta}_{*})}{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})}\,$, (17)
@@
 $\text{ with }\mathbf{LL}^{T}=\mathbf{\Sigma}=\frac{1}{S}\sum_{s=1}^{S}(\theta^{(s)}_{*}-\bar{\theta}_{*})(\theta^{(s)}_{*}-\bar{\theta}_{*})^{T}\,$, (19)
$\tilde{\mathbf{L}}\tilde{\mathbf{L}}^{T}=\frac{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})(\theta^{(s)}_{*}-\tilde{\theta}_{*})(\theta^{(s)}_{*}-\tilde{\theta}_{*})^{T}}{\sum_{s=1}^{S}w_{i}(\theta^{(s)}_{*})}\,.$ (20)
@@
Although IWMM is computationally more expensive than PSIS, we employ it in cases where PSIS fails to provide reliable approximations of the target distribution. Figure 2 illustrates representative posterior distributions from two imputed datasets in the simulation study described in Section 3.1, within the special case of the missing data problem (see Section 2.4). The proposal distribution (shown in red) was obtained via MCMC, and the true target distribution (yellow) — which is typically unknown without running MCMC — serves as the reference. The PSIS-based approximation (light blue), constructed via importance resampling from the proposal, fails to accurately capture the shape and location of the target distribution. This poor approximation is flagged by a high Pareto shape parameter, $\hat{k}=0.93>k_{\mathrm{threshold}}$}, indicating that the results are unreliable. In contrast, using additionally IWMM results in a substantially improved approximation (dark blue), with resampled draws that closely match the target distribution. This is confirmed by a much lower shape parameter, $\hat{k}_{\mathrm{MM}}=0.27<k_{\mathrm{threshold}}$}, indicating a successful and trustworthy approximation.
@@
We then define the mixture proposal as
@@
$r_{i}(\theta^{(s)}_{*})$ $=\frac{p(\theta^{(s)}_{*}\mid\tau^{(i)})}{p_{\rm mix}(\theta_{*}^{(s)}\mid\{\tau^{(j)}\})}$ (22)
$=\frac{p(\theta^{(s)}_{*}\mid\tau^{(i)})}{\frac{1}{\mid I_{\text{mix}}\mid}\sum_{j\in I_{\text{mix}}}p(\theta_{*}^{(s)}\mid\tau^{(j)})}$ (23)
@@
In the standard setting of self-normalized importance sampling, the normalizing constants — here the marginal likelihoods $p(\tau^{(i)})$ — can be ignored since the self-normalization automatically corrects for them. However, in the case of a mixture proposal, the constants of the mixture components cannot be factored out of the summation in the denominator, and thus self-normalization is no longer directly applicable. Therefore, to evaluate Equation (26), we require either the exact values or relative magnitudes of the marginal likelihoods $p(\tau^{(j)})$ for all $j\in I_{\text{mix}}$. For the purpose of estimating $p(\tau^{(j)})$, we use bridge sampling [50], a method for estimating normalizing constants in complex posterior distributions [30]. With these estimates, the mixture-based importance weights can be accurately computed, even though it adds computational complexity.
In our approach, we employ both single and mixture proposal distributions for PSIS, whereas IWMM is applied only with single proposals. Preliminary investigations (not shown here) indicated that using these mixtures within IWMM performs worse than using just a single posterior $p(\theta\mid\tau^{*})$ as proposal. We attribute this to the nature of the IWMM method, which relies on affine transformations to shift and scale the proposal distribution to match the moments of the weight distribution [56]. A single posterior distribution is more likely to share structural similarities — differing primarily in location and scale — with another posterior (the target) than a multi-modal mixture distribution. Based on these findings, we opted not to further consider mixture proposals when applying IWMM.